# An Survey of the Evolution of Language Model-Based Dialogue Systems

### v1

@@ 5.3 Internal Reasoning with External Sources
By focusing on enhancing the internal status of the dialogue context and allowing seamless interactions with external sources, LLM-based systems can become more sophisticated, effective, and capable of delivering personalized, trustworthy, helpful, up-to-date responses to various user queries and requests [234]. Recently, Tu et al. [210] utilize persona (internal) and profile memory (external) simultaneously to generate personalized responses while Zhang et al. [235] alternatively take advantage of short-term memory (internal) and long-term memory (external). Within the multi-persona collaboration framework TPE (Thinker, Planner, and Executor), Wang et al. [236] leverage the internal status of dialogue context to guide the planning use of different external knowledge sources, regarding different sources as one specific type of conceptual tools. In addition, Sumers et al. [191] propose CoALA, which delineates a language agent that incorporates modular memory components, a structured action space for interactions with both internal memory and external environments, and a generalized decision-making process for action selection.
6 Discussions
6.1 Multi-modality
@@
In the ever-evolving landscape of conversational AI, the integration of multimodal capabilities has emerged as a pivotal advancement that promises to reshape the future of LLM-based dialogue systems [237, 238, 239]. Before the advent of large language models, dialogue systems primarily operated within the confines of text-based interactions, with few works explore the multimodal dialogue systems [240, 241, 242, 243]. However, with the introduction of powerful models such as GPT-4 [177] and their ability to comprehend and generate text and even image at a human-like level, the stage was set for a transformative convergence of text, images, audio, and video in the realm of conversation [237, 238]. In this section, we first briefly review the previous work before the era of large langauge models and then explore the exciting prospects and future directions of multimodal LLM-based dialogue systems, shedding light on their evolution, potential applications, and the dynamic impact they promise to have on a wide range of fields, from customer service and medicine and beyond.
6.1.1 Before the era of LLMs
@@
Before the era of large language models, there were relatively few efforts which try to fuse different modalities seamlessly in the conversational flow. According to different modalities of input and output in dialogue systems, these works can be divided into two categories [^11] https://github.com/ImKeTT/Awesome-Multi-Modal-Dialog: 1) systems with diverse modalities in the input but not in the output [244, 245, 246, 247, 248]; 2) systems with multiple modalities in the output [249, 250, 251, 252]  [^12] This includes systems with the incorporation of multiple modalities in both the input and output.. Predominantly, a significant portion of research endeavors align with the first category, where input comprises text and image modalities, resulting in textual responses pertaining to the image content. A noteworthy milestone in this progression was achieved by Alamri et al. [245] when they pioneered the integration of audio information into the input phase, marking a notable enhancement in the multimodal dialogue system’s capabilities. It is essential to highlight that the majority of efforts falling under the second category primarily involve the generation of output in alternative modalities, such as images, employing a discriminative approach by selecting one from existing sources rather than generating one [240, 243]. Furthermore, some works explore multilingual situations in the context of multimodal dialogue [253, 254].
6.1.2 After the era of LLMs
@@
Since the emergence of LLMs, a notable trend has been the alignment of feature spaces across different modalities within dialogue systems. This alignment effort has taken two distinct directions: 1) the alignment of non-textual modalities with the textual feature space [237, 238, 255, 256, 257]; and 2) approaches that pursue alignment in the opposite direction [258, 259]. These endeavors mark a significant shift in the field, aiming to harmonize the representation and understanding of diverse modalities within the context of LLM-based dialogue systems. In detail, LLaVA represents the inaugural endeavor in visual instruction tuning using data generated by language-only GPT-4, serving as an end-to-end trained large multimodal model that connects a image encoder and LLM for general-purpose visual and language understanding [237]. Despite some preliminary efforts have been made in the development of image-based dialogue systems, some works specifically delve into the relatively uncharted territory of video-based dialogue systems [257, 256]. For example, Muhammad et al. [257] introduce Video-ChatGPT, which is a multimodal model that merges a video-adapted visual encoder with a LLM by simply finetuning a linear layer. Besides that, Zhang et al. [256] further present Video-LLaMA which is capable to understand both visual and auditory content in the video.
@@ 6.2 Hallucination
Language model-based dialogue systems have demonstrated significant potential across a broad range of applications. However, an increasing concern is their observed inclination towards hallucinations [6, 260]. These instances, where system-generated responses defy universally accepted word knowledge, present a substantial challenge to the reliability and effectiveness of these systems in real-world scenarios. As the utilization of these dialogue systems continues to evolve, understanding and mitigating the impact of these hallucinations becomes a critical task for their successful deployment.
6.2.1 Understanding the Causes of Hallucinations
@@
Understanding the underlying causes of hallucinations in language model-based dialogue systems is a crucial first step toward devising effective mitigation strategies. The hallucination phenomenon often results from a complex interplay of several factors. Inadequate training data [261, 262, 263] can play a significant role—when the training data lacks diversity or doesn’t cover certain topics, the model might attempt to fill in the gaps in its knowledge, leading to hallucinated responses. Model overgeneralization [264, 130] is another major contributor, where an language model might overgeneralize from the patterns it has learned during training, leading to responses that may seem plausible but are actually incorrect or nonsensical in a given context. Lastly, the inherent randomness in the generation process can also lead to hallucinations. The stochastic nature of language models [265, 266] means they can sometimes generate incorrect text, particularly when the model is unsure about the most appropriate response. Understanding these underlying causes provides a foundation to explore potential mitigation strategies.
6.2.2 Detecting Hallucinations in Dialogue System
@@
Research on hallucination detection has evolved along four main trajectories. Firstly, human evaluations have been conducted to assess the model output from diverse perspectives [267, 268]. However, these evaluations, while valuable, encounter issues due to their time-consuming nature and the inherent subjectivity of human judgement, leading to scalability and reproducibility challenges. Secondly, datasets have been constructed to evaluate the actuality of open-domain generation with the aid of references [266, 269]. These methods, while offering a more objective evaluation approach, are limited by their dependence on human-labelled references, which can negatively impact their applicability in real-world scenarios and their generalizability to dynamically generated content. Thirdly, self-evaluation methods [195, 270] have been developed to estimate a model’s uncertainty in its output. Despite the simplicity and appeal of these methods, they often lack interpretability and are less effective when dealing with long-form responses. Fourth, some recent studies [261, 271] have applied fact-checking principles to spot factual inaccuracies in the model’s output. These efforts, however, are often constrained by the limitations of external knowledge bases used in the fact-checking process. Finally, an emergent line of research focuses on integrating references or fingerprints into the generated content [272, 273]. This approach, while promising, introduces a layer of traceability and verifiability to the output, enabling users or regulators to trace and verify the origins and accuracy of the content.
6.2.3 Strategies for Mitigating Hallucinations

### v2

@@ 5.3. Internal Reasoning with External Acting
By focusing on enhancing the cognitive understanding of the dialogue context and allowing seamless interactions with external sources, LLM-based systems can become more sophisticated, effective, and capable of delivering personalized, trustworthy, helpful, up-to-date responses to various user queries and requests (Wang et al., 2023e; Acikgoz et al., 2025). For examples, Tu et al. (Tu and et al., 2023) utilize persona (internal) and profile memory (external) simultaneously to generate personalized responses while Xu et al. (Xu et al., 2022a) alternatively take advantage of short-term memory (i.e., internal reasoning) and long-term memory (i.e., external source). Within the multi-persona collaboration framework TPE (Thinker, Planner, and Executor), Wang et al. (Wang and et al., 2023d) leverage the internal status of dialogue context to guide the planning use of different external knowledge sources, regarding different sources as one specific type of conceptual tools. In addition, Sumers et al. (Sumers et al., 2023) propose CoALA, which delineates a language agent that incorporates modular memory components, a structured action space for interactions with both internal memory and external environments, and a generalized decision-making process for action selection.
Table 9. Performance of LLM-based Dialogue System which blender internal reasoning and external acting in an unified framework. The results are copied from CoALM  (Acikgoz and et al., 2025a).
Type Method MultiWoZ 2.4 API-Bank Success JGA Rouge-L BLEU-4 Trad. DS FNC-ToD 44.4 37.9 3.9 3.3 NC-Latent-ToD 68.3 39.7 3.2 0.8 LLM-Prompting Llama 3.1 8B Instruct 19.9 26.3 72.7 62.3 Granite-20B-Code 10.7 21.8 60.3 43.8 GPT4o-mini 69.9 38.4 - - LLM-Finetuning Hammer 2.0 7B 23.5 21.7 90.1 85.4 ToolAce 18.0 34.4 81.5 76.1 CoALM (8B) 51.6 30.4 92.8 89.4 CoALM (70B) 69.4 43.8 92.7 89.5
Despite these encouraging advancement to bridge the internal reasoning and external interaction, the better unification remains an open challenge due to the complex relation between them. Sometimes their functionalities may overlap with each other when one problem can be solve by both internal reasoning or external acting. Wang et al. (Wang and et al., 2025) propose Self-DC which unify reasoning and acting as different functions and dynamically invoke them based on self-aware knowledge boundary of LLMs while SMART (Qian and et al., 2025) further encompass a wider range of tool utilization scenarios, with a particular focus on addressing the issue of tool overuse for existing language agents. Besides that, there also exist conflict knowledge or results between internal reasoning and external acting (Xu and et al., 2024b), i.e., the conflict between the parametric knowledge and retrieved knowledge. More recently, Emre et al. (Acikgoz and et al., 2025b) propose a concept of conversational agents, where an LLM-based dialogue system is designed to perform multi-turn interactions with users, by integrating reasoning and planning capabilities with action execution. Combining the advantages of both language agent and dialogue system, conversational agent is capable to master both multi-turn conversation and tool use with an unified model (Acikgoz and et al., 2025a). Table 9 shows the results between CoALM with other strong baselines. It can be drawn that CoALM achieves leading performance on both TOD and LA benchmarks, revealing the strong robustness and generalization. In conclusion, ensuring seamless integration between internal reasoning and external acting enables better monitoring and controllability (Wang et al., 2025c). Addressing these issues is essential for developing more reliable, adaptive, and intelligent LLM-based dialogue systems capable of engaging in complex and dynamic conversations.
5.4. Summary
Throughout the evolution of dialogue systems, we have consistently observed a clear trend: the pursuit of a powerful, scalable, and unified system. From the independent development of TOD and ODD to their gradual fusion at various levels, and now the rise of LLM-based dialogue systems, the ultimate goal remains the same. As a consequence, existing LLM-based dialogue systems enables more natural, personalized, and trustworthy interactions that adapt to complex real-world scenarios, as exemplified by architectures like TPE (Wang and et al., 2023d) and CoALA (Sumers et al., 2023). However, significant challenges persist. First, ethical concerns emerge regarding privacy preservation when handling sensitive user profiles (§ 5.2) and psychological manipulation risks in emotionally-aware systems (§ 5.1). Second, one notable drawback is the issue of hallucinations, where the system may generate inaccurate or misleading information. Despite external knowledge source may alleviate the hallucinations in the response, it introduces additional hallucination issues in the planning step of dialogue system. Overall, while LLM-based dialogue systems have pushed the boundaries of what traditional dialogue system can achieve, further research is essential to ensure their reliability, ethical compliance, and practical deployment in complex environments.
6. Discussions
6.1. Data
As large language model (LLM)-based dialogue systems continue to evolve towards becoming integral components of everyday life and potential gateways to artificial general intelligence (AGI), the discussion around data becomes increasingly pivotal. Two central aspects of data issues emerge: general challenges related to data schema and quality, and specific challenges concerning diversity in terms of languages, domains and applications.
6.1.1. General Data Issues: Universal Data Schema and Quality
One key observation we can drawn from previous evolution of dialogue system is the unification of data schema or format is the very foundation which enables the the unification of tasks and models. As the LLM-based dialogue system become increasingly embedded in real-world applications to solve more and more complex task, it proposes new challenges about new data schema and format to learn and complete these new tasks together. On the schema side, several pioneer attempt to unify all tasks into one unified data format. For example, CoALM (Acikgoz and et al., 2025a) adopts ReAct-style data schema to empower the LLM-based dialogue system capable to function as both dialogue system and language agent, aiming to unify multi-turn conversation, function calling and task completion. On the quality side, it is difficult and expensive to manually collect the required high-quality data for each task and application, and thus most of researcher try to utilize LLMs themselves to synthesize data (Wang and et al., 2023f), leading in self-evolving LLMs (Yuan and et al., 2024; Tao and et al., 2024). For examples, Yuan et al. (Yuan and et al., 2024) study Self-Rewarding Language Models, where the language model itself is used via LLM-as-a-Judge prompting to provide its own rewards during training, achieving iterative improvement by finetuning using self-rewarding self-generated responses, followed by lots of works (Huang and et al., 2025; Hongru and et al., 2025). Fruthermore, as the model approaches or surpasses human-level intelligence, its performance may become limited by the quality and availability of human-labeled data. This creates a challenge in scaling data acquisition, such as weak-to-strong supervision (Burns and et al., 2023), making it an open question that remains to be addressed.
6.1.2. Specific Data Issues: Diversity in Language, Application, and Domain
While the above general issues affect the performance of LLM-based dialogue systems, specific challenges arise when addressing diverse, low-resource languages and various application domains. Although we previously discussed existing work on low-resource languages, a significant gap remains in resources and applications compared to high-resource languages. This imbalance has far-reaching consequences, as dialogue systems may struggle to deliver equitable and culturally aware services, ultimately reinforcing existing inequalities and cultural biases. Addressing these challenges involves collecting and curating data that reflects the linguistic, cultural, and domain-specific diversity of real-world interactions. Additionally, dialogue systems designed for niche domains – such as medical advice, legal consultations, or finance – demand domain-specific data that accurately captures the specialized vocabulary, contextual nuances, and ethical considerations inherent to those fields. By emphasizing scalability and inclusivity in data collection and processing, dialogue systems can become more resilient, equitable, and better equipped to meet the diverse and complex needs of our interconnected world.
6.2. Applications
This section examines real-world applications of dialogue systems, categorized into three types based on use cases: Information-Seeking and Decision Support, Task Orchestration and Execution, and Affective and Recreational Engagement. These systems are deployed using either standalone architectures, which operate independently, or integrated architectures, which connect to external tools, knowledge bases, or other components. We classify each system accordingly and evaluate their relative advantages. A key trend across all categories is the growing effectiveness of LLM-based dialogue systems when integrated with external resources. Recent systems like GPT-4 (OpenAI, [n. d.]), Claude  [^14] https://www.anthropic.com/news/claude-3-7-sonnet, and DeepSeek (DeepSeek-AI, 2025) exemplify this, functioning as general-purpose agents enhanced by access to specialized tools. Despite the rise of such versatile systems, specialized dialogue applications remain vital for scenarios with constraints such as limited compute, domain-specific needs, or strict privacy requirements.
6.2.1. Information-seeking and Decision-support.
This type of applications help users access information and make informed decisions through conversational interfaces. We introduce three representative applications: 1) E-commerce Customer Support. Dating back to the 1990s, Meng and et al. (1996) introduced WHEELS, an Integrated dialogue system using parsing strategy for NLU and form-filling for dialogue management, which functioned as a natural language interface to an automobile advertisement database. With the rise of deep learning, leading e-commerce companies adopted Standalone models trained on large-scale corpus (Chen and et al., 2020). Latest implementations integrate product knowledge bases and multimodal interactions, enhancing customer experience through personalized services while maintaining conversational coherence (Gupta and et al., 2024); 2) E-government Civil Service. Recently, government agencies worldwide have widely adopted dialogue systems to help citizens access information and navigate services (Watch, 2022; Abed, 2024). For instance, Singapore’s GovTech team presents “Ask Jamie” [^15] https://www.tech.gov.sg/media/technews/govtech-team-behind-ask-jamie-government-chatbot/, a chatbot to address citizen inquires for government agencies. Initially deployed as a standalone system trained on extensive QA documents, latest government implementations now leverage integrated approaches using RAG techniques with capable dialogue systems as the backbone [^16] https://insidegovuk.blog.gov.uk/2024/01/18/the-findings-of-our-first-generative-ai-experiment-gov-uk-chat/, effectively addressing data privacy concerns and mitigating hallucination issues; 3) Financial Robo-advisors. Adopting dialogue systems in finance to provide asset and wealth management advice has gained significant traction with the advent of LLM-based solutions (Polireddi, 2024). In this sector, Integrated systems that combine LLMs with domain-specific techniques to deliver sophisticated advisory services are prevalent. Leading companies like Wealthfront have incorporated dialogue systems into their automated investment platforms [^17] https://research.wealthfront.com/whitepapers/investment-methodology/, leveraging seamless interactive capabilities to deliver personalized financial guidance and optimize portfolio management at substantially lower costs than traditional human advisors. Beyond the previously mentioned fields, dialogue systems are also widely adopted in domains requiring specialized expertise including: medical advice systems (e.g., Google’s Med-PaLM (Singhal and et al., 2025)), educational assistants (e.g., Khanmigo [^18] https://www.khanmigo.ai/), and legal question-answering services (Yao and et al., 2024).
6.2.2. Task Orchestration and Execution
Dialogue systems is widely adopted to help users accomplish specific goals through conversational interfaces, ranging from simple standalone helpers to complex integrated systems that coordinate multiple tools and services. We review such applications in following two categories: 1) Personal Assistance. Dialogue systems functioning as personal assistants are usually embedded in diverse smart devices, requiring tight integration between natural language processing capabilities and hardware interfaces or APIs. Early implementations like Siri and Alexa provides relatively basic command execution, while latest LLM-based systems such as Claude and ChatGPT leverage more sophisticated semantic understanding and reasoning to deliver sophisticated assistance through advanced tool integration for personal task management; 2) Specialized Expert Assistance. Different from aforementioned personal assistants, such expert assistance DS needs to execute professional tasks with domain expertise. For example, DoNotPay  [^19] https://donotpay.com/ functions as a legal assistant that generates properly formatted legal documents and navigates procedural requirements through conversational interfaces. BloombergGPT (Wu and et al., 2023b) operates as a pre-trained standalone LLM optimized for finance-specific tasks like news sentiment analysis. FinRobot (Yang and et al., 2024) implements an integrated architecture where the LLM serves as the “brain”, i.e., a central orchestration module, for a financial agent, coordinating specialized tools for financial analysis, market prediction, etc. More recently, LLM-based DS have unlocked new frontiers in various fields, enabling autonomous problem-solving and complex decision-making in real-world tasks, such as ChatDEV (Qian and et al., 2023) for software development, Copilot [^20] https://copilot.microsoft.com/ for coding assistance, and GUI agents (Wang and et al., 2024) that execute interface actions based on user instructions.
6.2.3. Affective and Recreational Engagement
Dialogue systems for affective engagement serve dual purposes: facilitating recreational social interactions and providing emotional support through therapeutic conversations. Thus they operate on a spectrum from casual companionship to clinical intervention. Commercial implementations like Replika  [^21] https://replika.com/ primarily function as Standalone systems offering companionship through social chit-chat, maintaining substantial user engagement through simulated emotional bonds. They also show clinical benefits, with research confirming their effectiveness in reducing suicidal thoughts through timely emotional support and crisis intervention. More specialized therapeutic systems usually integrate external tools to provide more professional service (Filienko and et al., 2024). WoeBot Health  [^22] https://woebothealth.com/ utilizes persistent databases to track user emotional states and deliver personalized mental health interventions, while Wysa  [^23] https://www.wysa.com/ employs a hybrid approach connecting automatic conversations with emergency intervention APIs and human therapist escalation when needed.
6.3. Multi-modality
@@
In the ever-evolving landscape of conversational AI, the integration of multimodal capabilities has emerged as a pivotal advancement that promises to reshape the future of LLM-based dialogue systems (Liu et al., 2023a; Zheng and et al., 2023). Before the advent of large language models, dialogue systems primarily operated within the confines of text-based interactions, with few works explore the multimodal dialogue systems (Johnston et al., 2002; Liu et al., 2022a; Liu and et al., 2022; Liao et al., 2018). However, with the introduction of powerful models such as GPT-4 (OpenAI, [n. d.]) and their ability to comprehend and generate text and even image at a human-like level, the stage was set for a transformative convergence of text, images, audio, and video in the realm of conversation (Liu et al., 2023a). In this section, we first briefly review the previous work before the era of large langauge models and then explore the exciting prospects and future directions of multimodal LLM-based dialogue systems, shedding light on their evolution, potential applications, and the dynamic impact they promise to have on a wide range of fields, from customer service and medicine and beyond.
6.3.1. Before the era of LLMs
@@
Before the era of large language models, there were relatively few efforts which try to fuse different modalities seamlessly in the conversational flow. According to different modalities of input and output in dialogue systems, these works can be divided into two categories [^24] https://github.com/ImKeTT/Awesome-Multi-Modal-Dialog: 1) systems with diverse modalities in the input but not in the output (Das et al., 2017; AlAmri et al., 2019; Shuster and et al., 2020; Wang and et al., 2021b; Zheng and et al., 2022); 2) systems with multiple modalities in the output (Saha et al., 2018; Zang and et al., 2021; Ma and et al., 2022; Feng and et al., 2023)  [^25] This includes systems with the incorporation of multiple modalities in both the input and output.. Predominantly, a significant portion of research endeavors align with the first category, where input comprises text and image modalities, resulting in textual responses pertaining to the image content. A noteworthy milestone in this progression was achieved by Alamri et al. (AlAmri et al., 2019) when they pioneered the integration of audio information into the input phase, marking a notable enhancement in the multimodal dialogue system’s capabilities. It is essential to highlight that the majority of efforts falling under the second category primarily involve the generation of output in alternative modalities, such as images, employing a discriminative approach by selecting one from existing sources rather than generating one (Johnston et al., 2002; Liao et al., 2018). Furthermore, some works explore multilingual situations in the context of multimodal dialogue (Hough and et al., 2016).
6.3.2. After the era of LLMs
@@
Since the emergence of LLMs, a notable trend has been the alignment of feature spaces across different modalities within dialogue systems. This alignment effort has taken two distinct directions: 1) the alignment of non-textual modalities with the textual feature space (Liu et al., 2023a; Maaz and et al., 2023); and 2) approaches that pursue alignment in the opposite direction (Girdhar et al., 2023). These endeavors mark a significant shift in the field, aiming to harmonize the representation and understanding of diverse modalities within the context of LLM-based dialogue systems. In detail, LLaVA represents the inaugural endeavor in visual instruction tuning using data generated by language-only GPT-4, serving as an end-to-end trained large multimodal model that connects a image encoder and LLM for general-purpose visual and language understanding (Liu et al., 2023a). Despite some preliminary efforts have been made in the development of image-based dialogue systems, some works specifically delve into the relatively uncharted territory of video-based dialogue systems (Maaz and et al., 2023). For example, Muhammad et al. (Maaz and et al., 2023) introduce Video-ChatGPT, which is a multimodal model that merges a video-adapted visual encoder with a LLM by simply finetuning a linear layer.
@@ 6.4. Hallucination
Language model-based dialogue systems have demonstrated significant potential across a broad range of applications. However, an increasing concern is their observed inclination towards hallucinations (Huang et al., 2021; Ji et al., 2023). These instances, where system-generated responses defy universally accepted word knowledge, present a substantial challenge to the reliability and effectiveness of these systems in real-world scenarios. As the utilization of these dialogue systems continues to evolve, understanding and mitigating the impact of these hallucinations becomes a critical task for their successful deployment.
6.4.1. Understanding the Causes of Hallucinations
@@
Understanding the underlying causes of hallucinations in language model-based dialogue systems is a crucial first step toward devising effective mitigation strategies. The hallucination phenomenon often results from a complex interplay of several factors. Inadequate training data (Kandpal et al., 2023; Xue and et al., 2023) can play a significant role—when the training data lacks diversity or doesn’t cover certain topics, the model might attempt to fill in the gaps in its knowledge, leading to hallucinated responses. Model overgeneralization (Hupkes et al., 2020; Lin et al., 2022) is another major contributor, where an language model might overgeneralize from the patterns it has learned during training, leading to responses that may seem plausible but are actually incorrect or nonsensical in a given context. Lastly, the inherent randomness in the generation process can also lead to hallucinations. The stochastic nature of language models (Ranzato et al., 2016; Lee et al., 2023) means they can sometimes generate incorrect text, particularly when the model is unsure about the most appropriate response. Understanding these underlying causes provides a foundation to explore potential mitigation strategies.
6.4.2. Detecting Hallucinations in Dialogue System
@@
Research on hallucination detection has evolved along four main trajectories. Firstly, human evaluations have been conducted to assess the model output from diverse perspectives (Yu et al., 2023; Liu et al., 2023c). However, these evaluations, while valuable, encounter issues due to their time-consuming nature and the inherent subjectivity of human judgement, leading to scalability and reproducibility challenges. Secondly, datasets have been constructed to evaluate the actuality of open-domain generation with the aid of references (Lee et al., 2023). These methods, while offering a more objective evaluation approach, are limited by their dependence on human-labelled references, which can negatively impact their applicability in real-world scenarios and their generalizability to dynamically generated content. Thirdly, self-evaluation methods (Varshney and et al., 2023) have been developed to estimate a model’s uncertainty in its output. Despite the simplicity and appeal of these methods, they often lack interpretability and are less effective when dealing with long-form responses. Fourth, some recent studies (Min and et al., 2023) have applied fact-checking principles to spot factual inaccuracies in the model’s output. These efforts, however, are often constrained by the limitations of external knowledge bases used in the fact-checking process. Finally, an emergent line of research focuses on integrating references or fingerprints into the generated content (Gao et al., 2023). This approach, while promising, introduces a layer of traceability and verifiability to the output, enabling users or regulators to trace and verify the origins and accuracy of the content.
6.4.3. Strategies for Mitigating Hallucinations